# Bi-VLA: Vision-Language-Action Model-Based System for Bimanual Robotic Dexterous Manipulations

### v1

@@ Abstract
This research introduces the Bi-VLA (Vision-Language-Action) model, a novel system designed for bimanual robotic dexterous manipulations that seamlessly integrate vision, language understanding, and physical action. The system’s functionality was evaluated through a set of household tasks, including the preparation of a desired salad upon human request. Bi-VLA demonstrates the ability to interpret complex human instructions, perceive and understand the visual context of ingredients, and execute precise bimanual actions to assemble the requested salad. Through a series of experiments, we evaluate the system’s performance in terms of accuracy, efficiency, and adaptability to various salad recipes and human preferences. Our results indicate a high success rate of 100% in generating the correct executable code by the Language module from the user-requested tasks. The Vision Module achieved a success rate of 96.06% in detecting specific ingredients and an 83.4% success rate in detecting a list of multiple ingredients.
@@ I Introduction
Recent advancements in language models have had a significant impact on Human-Robot Interaction (HRI), enabling robots to interact more naturally and effectively with their human counterparts [1]. To enhance performance efficiency and deal with complex tasks, coordinated dual-arm robotic systems are used to increase reliability and approach tasks that are hard to be done with one manipulator [2]. The coordination of two manipulator systems has been studied in the literature to achieve satisfactory performance [3],[4]. However, traditional planning-based methods that focus on motion planning for coordinating two robotic arms through predefined trajectories [5], [6] are not optimal to be applied for dynamic and complex tasks. To achieve human-level bimanual manipulation, strategies like reinforcement learning (RL) [7] and learning from demonstration (LfD) [8] can be used. However, it’s crucial to acknowledge the inherent challenges associated with these methods, particularly concerning the time required for learning and data collection. Moreover, they lack robustness in dynamically changing environments and tasks, limiting their practical utility in scenarios requiring real-time responsiveness to human needs.
@@
Using Language models in HRI not only facilitates natural communication but also extends their utility beyond basic interactions, encompassing tasks such as task planning and generating intermediate reasoning steps, commonly referred to as a chain of thought. However, the use of Language models to synthesize the bimanual skills of robots has not received significant attention yet, although numerous tasks that humans perform routinely require the coordinated use of both hands, emphasizing the crucial role of bimanual capabilities in executing intricate operations beyond the reach of single-arm manipulators.
Expanding on this premise, the integration of vision, language, and action in a unified model presents a novel approach to enhancing robotic bimanual operation. In this research, we aim at the development of a vision-language-action model, dubbed as Bi-VLA, for bimanual robotic operation. We also aim to explore its potential impact in orchestrating the coordination of dual-arm robotic systems in response to human instructions.
Figure 2: The LLM-based semantic planner receives user instructions and generates a detailed plan outlining the movements of two robot manipulators. The generated plan is translated into a set of motion API call functions inside a Python function through the LLM-based Code Generator. The execution of the generated code triggers the movement of the two robot manipulators through the Motion Controller.
@@ II Related Work
In this section, we explore previous studies that examine how language instructions can be used to generate robot actions, along with prior research on bimanual robotic manipulation. Additionally, we offer insights into how this work contributes to the existing literature.
@@ II-A From language instructions to robotic actions
Translating user language input into robotic action has seen considerable progress through the integration of Large Language Models (LLMs) and various algorithmic frameworks. CLIPort [9] leverages visual and language models for robotic manipulation, focusing on the spatial understanding necessary for accomplishing ’where’ tasks. In a similar vein, PROGPROMPT [10] has innovated in generating situated robot task plans, utilizing LLMs for defining and scoring potential actions without the need for explicit domain knowledge. The "Code as Policies"[11] paradigm further underscores the importance of programmatic structures in embodying LLM-generated instructions for control, which is in close alignment with RT-1[12]’s objective of scalable, real-world robotic control. Drawing upon pre-trained vision-language models, open-world object manipulation [13] is evolving to adapt robotic actions to diverse and changing environments. Instruct2Act [14] steps beyond basic instructions by aligning multimodal inputs with actionable robotic sequences, aided by the breadth of responses offered by LLMs. "Language to Rewards" [15] posits an intriguing method of reward function generation from natural language inputs to calibrate and optimize robotic skills, while the goal of building cooperative agents finds promise in harnessing modularity through LLM architectures. Continuing this thread, RT-2 [16] focuses on the translation of visual perceptions and user language into actionable instructions, demonstrating the critical intersection between understanding and physical task execution. Lastly, PaLM-E signifies the convergence of multimodal understanding within LLMs, emphasizing the embodiment of language models in tangible, robotic actions. Collectively, these landmarks in research contribute to an overarching framework that is not only advancing the potential of robots to understand and act upon complex user language inputs but also redefining interactions between humans and robots. In this work, we use LLMs as a strategic semantic planner, orchestrating the collaborative efforts of the robotic arms to successfully complete the given task.
@@ II-B Bimanual robotic manipulation
The exploration of bimanual object manipulation spans a multidisciplinary field involving insights from neuroscience, robotics, and computer science. In establishing a foundation, researchers have proposed various taxonomies for bimanual manipulation, acknowledging the unparalleled human capability for coordinating hands in complex tasks that robotics strives to emulate. The integration of spatial and temporal constraints is pivotal in achieving such dexterity and refined manipulative actions [17]. Extending from taxonomy, behavioral control paradigms suggest that leveraging dual-arm coordination can outperform single-arm approaches in manipulation tasks, emphasizing the significance of constructing behavior-based control systems for more adept bimanual robotic interaction [18]. A breadth of surveys and studies collectively push the frontier of dual-arm manipulation further, considering constraint-based programming as a promising avenue for dictating robotic precision and adaptability in varied manipulation scenarios [19]. In particular, haptic feedback stands out as a cornerstone for enhancing bimanual robotic systems, aiming to replicate the human sense of touch and force dynamics, thus empowering robots to manipulate objects with the nuanced force and feedback necessary in delicate or complex operations [20].
The evolution of bimanual robotic manipulation continues to address the ubiquitous uncertainties present in real-world interactions. Models and frameworks have been developed to delineate strategies for dealing with unpredictability in object properties and environmental conditions, fortifying the robustness of dual-arm manipulation under such challenges [21]. Crucially, the concept of shared control blurs the lines between human and robot manipulation, allowing for a symbiosis that can capitalize on the strengths of both parties in accomplishing bimanual tasks [4]. Further extending the accessibility and practicality of bimanual manipulation, research delving into low-cost teleoperation solutions like Mobile ALOHA[22] has opened avenues for learning complex manipulation behaviors with economical means and also unfolds in studies that leverage deep learning for bimanual robotic manipulation, illustrating an imitation learning framework conducive to teaching robots through observation and replication of human actions [18]. Progress in this area is complemented by advancements in techniques such as sim-to-real deep reinforcement learning, which fine-tunes bimanual tactile manipulation abilities in simulated environments before transferring them to real-world applications [23]. In contrast to the majority of the methods mentioned earlier, our research centers on a less data-hungry approach that doesn’t necessitate training or additional fine-tuning. In this method, the language module is used to call specific predefined action APIs and determine their calling sequence, along with allocating tasks between the two robot arms to complete the user’s assigned task.
@@ III Grounding Language Instructions into Bimanual Robot Actions
In this section, we offer insights into the architecture and data flow between the language, visual, and actions APIs modules of our proposed approach.
III-A Large Language Model Module
III-A1 Semantic planning
Semantic planning involves the use of natural language understanding and generation techniques to interpret human instructions and generate executable plans. In the context of our work, the semantic planner is designed to generate a plan for coordinating the movements of the two robot arm manipulators equipped with different and appropriate tools in order to perform the given task. The planner takes a task description as input and produces a plan specifying the sequence of actions that need to be executed by the robot arms (Fig. 2 left).
The planner utilizes a template that provides a general structure for the plan. It starts with the "[start of plan]" tag and ends with the "[end of plan]" tag, ensuring that the plan is clearly delineated. The template also includes specific instructions and placeholders that need to be filled based on the task requirements by the vision language model (see section III-B1), as well as a set of rules to follow.
In this work, we used Starling-LM-7B-alpha [24], an open-source LLM trained by Reinforcement Learning from AI Feedback (RLAIF). Starling-LM-7B-alpha was developed by researchers at UC Berkeley’s NEST Lab and it has demonstrated impressive performance, surpassing GPT-3.5 [25] in several benchmarks, including MT Bench, where it scored 8.09 out of 10.
@@ III-A1 Semantic planning
III-A2 Translating semantic plan into APIs function calls with the LLM Based Code Generator
In the second stage, the plan generated by the semantic planner is translated into a sequence of API call functions. The goal is to convert the textual plan into executable Python code that adheres to a predefined set of functions provided by the API. We use a prompt that outlines the functions available for use, the structure of the plan, and the rules to follow. To illustrate the process, we provide an example plan and its corresponding Python code (Fig. 2 middle), helping to guide the model’s response as in the one-shot prompting technique. This methodology facilitates the seamless translation of high-level task plans into executable Python code, promoting efficiency and clarity.
III-A3 From API function call to Action Generation with The Motion Control APIs
@@
  • open_gripper(manipulator_type): Completely opens the gripper.
  • move_to_object(manipulator_type, object_name): Moves the manipulator to the position of a specified object.
  • grasp(manipulator_type, object_name): grasps an object.
Additionally, we have developed API functions to control the specific type of task to be performed by the manipulator arms:
  • cut(manipulator_type, object_name): For cutting tasks.
  • pour(manipulator_type, object_name): For pouring tasks.
  • put(manipulator_type, object_name): For pick-and-place tasks.
  • toss(manipulator_type, object_name): Used for mixing objects or vegetables in a bowl.
  • cut_and_put_in(manipulator_type, object_name): That combines the cut and put functions.
The vision-language model uses the function get_list_of_objects() to retrieve the names of all objects visible on the table within the camera frame, as well as get_bounding_boxes() which generates the bounding boxes of the desired list of objects. Ensuring explicit naming of API functions is crucial to preventing confusion with the language model. This comprehensive set of API functions provides the necessary control and versatility for robotic operations, enabling a wide range of complex tasks to be performed with precision and efficiency.
III-B Vision Language Model Module
The Vision Language Model Module consists of various interconnected components that seamlessly collaborate to fulfill its intended purpose. This section delves into all the components encompassed within the Module, as well as their interconnectivity and interactions with one another.
III-B1 Integration of the Qwen Vision-Language Model
The QWEN-VL series are large-scale vision-language models (VLMs) designed to perceive and understand both text and images [26]. They are built upon the QWEN-LM language model foundation and endowed with visual capabilities through a carefully designed architecture and training pipeline. The models demonstrate state-of-the-art performance on a wide range of vision-language benchmarks, including image captioning, visual question answering, and visual grounding. In this work, we use the model to serve as a tool that verifies the availability of all required items to accomplish the task, as well as providing the pixel coordinates of desired items.
III-B2 Mapping Image Pixel to 3D Coordinates
We assume that the position of an object is equal to the position of its center. The relationship between the position of an object in the world coordinate system and pixel coordinates is given by the equation:
@@ III-B2 Mapping Image Pixel to 3D Coordinates
${}^{c}_{w}\grave{T}=\left[\begin{array}[]{c|c}{}^{c}_{w}R_{3\times 3}&{}^{c}_{w}t_{3\times 1}\end{array}\right]$
@@
${}^{c}_{w}T=\left[\begin{array}[]{cccc}&{}^{c}_{w}\grave{T}&&\\ \hline\cr 0&0&0&1\end{array}\right]=\left[\begin{array}[]{c|c}{}^{c}_{w}R_{3\times 3}&{}^{c}_{w}t_{3\times 1}\\ \hline\cr 0_{1\times 3}&1\end{array}\right]$
@@
IV Experiments
IV-A Experiments Setup
Two collaborative robots UR3 from Universal Robots were mounted on a Siegmund welding table, as shown in Fig. 1. A camera holder was designed and 3D printed in PLA material and installed in the manipulator with a gripper, a Logitech HD 1080p webcam, and a two-finger Robotiq gripper. A knife holder was designed, 3D printed, and installed in the end effector of the manipulator with the tool. Close to the manipulator with the gripper, the ingredients to prepare the different salads were placed, and the camera, installed in its end-effector, captured a picture to grasp the required ingredients as required. Between the two manipulators, a cutting board was located, where the cutting operation was performed. In front of the cutting board, a plate/bowl was located, in which the ready ingredients were placed.
Figure 4: Step-by-step execution of the Semantic Plan: (a)-(b) the manipulator with the gripper moves to grasp the pepper and bring it to the cutting board; (c)-(d) the manipulator with the tool moves to the cutting board, cuts the pepper, and places it in the bowl; (e)-(f) the two manipulators return to their initial position.
@@ IV-B Experiments Workflow
As depicted in Fig. 3, we design several custom cooking recipes stored in a vector database, which serves as a set of tasks for the robot manipulators to perform collaboratively. The recipes clearly state the required ingredients as well as the instructions to follow for successful execution. Among these sets of tasks, we evaluate the performance of our system on three selected tasks, including making a Vegetable Salad, a Russian Salad, and a Fruit Salad. The three tasks are different in nature in the sense that they require several types of actions, such as picking, placing, cutting, pouring, mixing, etc.
When a user submits a request, the request message serves as a query that is processed through the LLM-Based Retrieval-Augmented Generation (RAG) technique. RAG is a method that aims to improve the quality and factuality of responses generated by large language models (LLMs) [29]. It achieves this by augmenting the model’s outputs with relevant information retrieved from external knowledge sources, in this case, a vector database. This helps to enhance the accuracy and factual grounding of the generated responses.
After the RAG model produces a response, the VLM module checks the availability of all the ingredients mentioned in the recipe. Based on the VLM’s response, the LLM module will then generate a set of steps or sub-tasks for the two robots to perform if the ingredients are available. Alternatively, if the ingredients are not available, the LLM module will send an instruction to the robots not to perform any further actions.

### v2

@@ Abstract
This research introduces the Bi-VLA (Vision-Language-Action) model, a novel system designed for bimanual robotic dexterous manipulation that seamlessly integrates vision for scene understanding, language comprehension for translating human instructions into executable code, and physical action generation. We evaluated the system’s functionality through a series of household tasks, including the preparation of a desired salad upon human request. Bi-VLA demonstrates the ability to interpret complex human instructions, perceive and understand the visual context of ingredients, and execute precise bimanual actions to prepare the requested salad. We assessed the system’s performance in terms of accuracy, efficiency, and adaptability to different salad recipes and human preferences through a series of experiments. Our results show a 100% success rate in generating the correct executable code by the Language Module, a 96.06% success rate in detecting specific ingredients by the Vision Module, and an overall success rate of 83.4% in correctly executing user-requested tasks.
Keywords— Vision-Language-Action Model, Bimanual Robotic Manipulation, Human-Robot Interaction, Generative AI.
@@ I Introduction
Recent advancements in language models have significantly impacted Human-Robot Interaction (HRI), enabling robots to engage more naturally and effectively with their human counterparts [1]. In particular, to enhance performance efficiency and tackle complex tasks, coordinated dual-arm robotic systems are increasingly employed. These systems offer greater reliability and facilitate the execution of tasks that are challenging for a single manipulator [2]. Furthermore, the coordination of two manipulator systems has been extensively studied in the literature, yielding satisfactory performance outcomes [3], [4]. However, traditional planning-based methods, which focus on motion planning through predefined trajectories for coordinating two robotic arms [5], [6], are often suboptimal for dynamic and complex tasks. To achieve human-level bimanual manipulation, alternative strategies such as reinforcement learning (RL) [7] and learning from demonstration (LfD) [8] have been proposed. Nevertheless, it is crucial to acknowledge the inherent challenges associated with these approaches, particularly regarding the time required for learning and data collection.
@@
Moreover, the integration of language models in HRI not only facilitates natural communication but also extends their utility beyond basic interactions. This includes tasks such as task planning and generating intermediate reasoning steps, commonly referred to as a chain of thought. However, despite their potential, the application of language models to synthesize the bimanual skills of robots has not received significant attention. This is particularly noteworthy given that numerous tasks humans perform routinely require the coordinated use of both hands, highlighting the essential role of bimanual capabilities in executing intricate operations that are beyond the reach of single-arm manipulators. Building on this premise, the integration of vision, language, and action within a unified model presents a novel approach to enhancing robotic bimanual operation. In this research, we aim to develop a vision-language-action model, dubbed Bi-VLA, specifically designed for bimanual robotic operations. Additionally, we will explore its potential impact on coordinating the actions of dual-arm robotic systems in response to human instructions. The proposed scenario is a cooking robot, where a robot is in charge of picking and placing the required ingredients, and the second cutting them.
Figure 2: Bi-VLA Architecture.
@@ II-A From language instructions to robotic actions
Translating user language input into robotic actions has seen considerable progress through the integration of Large Language Models (LLMs) and various algorithmic frameworks. CLIPort [9] leverages visual and language models for robotic manipulation, focusing on the spatial understanding necessary for accomplishing the tasks. Similarly, ProgPrompt [10] introduced a programmatic LLM prompt structure that enables plan generation for robotic tasks scoring potential actions without the need for explicit domain knowledge. The “Code as Policies" [11] paradigm further underscores the importance of programmatic structures in embodying LLM-generated instructions for control, which is in close alignment with RT-1’s objective of scalable, real-world robotic control [12]. Drawing on pre-trained vision-language models, open-world object manipulation [13] is evolving to adapt robotic actions to diverse and changing environments. Instruct2Act [14] steps beyond basic instructions by aligning multimodal inputs with actionable robotic sequences, aided by the breadth of responses offered by LLMs. “Language to Rewards" [15] introduced a method of reward function generation from natural language inputs to calibrate and optimize robotic skills, with the goal of building cooperative agents by harnessing modularity through LLM architectures. Continuing this thread, RT-2 [16] focuses on translating visual perceptions and user language into actionable instructions, demonstrating the critical intersection between understanding and physical task execution. Lastly, PaLM-E[17] highlights the convergence of multimodal understanding within LLMs, emphasizing the embodiment of language models in tangible robotic actions, while LLM-MARS[18] incorporates LLMs into the multi-agent robotics system for behavior trees generation. Collectively, these advancements contribute to an overarching framework that not only enhances robots’ ability to understand and act on complex user language inputs but also redefines interactions between humans and robots. In this work, we use LLMs as a strategic semantic planner, orchestrating the collaborative efforts of robotic arms to successfully complete the given task.
@@ II-B Bimanual robotic manipulation
The exploration of bimanual object manipulation spans a multidisciplinary field, incorporating insights from neuroscience, robotics, and computer science. To establish a foundation, researchers have proposed various taxonomies aimed at emulating human-like coordination in complex tasks. Notably, the integration of spatial and temporal constraints is essential for achieving dexterity [19]. Moreover, behavioral control paradigms suggest that dual-arm coordination often surpasses single-arm methods, highlighting the necessity for behavior-based control systems to facilitate effective bimanual interaction, as demonstrated by studies utilizing deep learning for imitation learning that illustrate effective teaching methods for robots through the observation and replication of human actions [20]. Additionally, low-cost teleoperation solutions, such as Mobile ALOHA [21], have opened avenues for learning complex manipulation behaviors. Finally, progress in this area is complemented by advancements in sim-to-real deep reinforcement learning, which refines bimanual tactile manipulation abilities in simulated environments before transferring them to real-world applications [22]. In this work, we focus on a less data-intensive approach that eliminates the need for training or additional fine-tuning. This method involves the language module invoking specific predefined action APIs, determining their calling sequence, and allocating tasks between the two robotic arms to complete the user’s assigned task.
@@ III Grounding Language Instructions into Bimanual Robot Actions
III-A General system overview
The system takes as input the user’s task request, which is transcribed into a text message, along with an image from the camera (see Fig. 2 left). The task request serves as a query for the Retrieval-Augmented Generation (RAG) system, and its output initiates a visual question-answering (VQA) dialogue between the Language Module and the Vision Language Module. The input image, on the other hand, is split into patches, tokenized, and combined with the VQA question tokens before being sent to the Vision Language Module. The main outputs include an image with captions and a JSON file containing the 2D positions of ingredients, a list of all ingredients, and missing ingredients, etc. Subsequently, the Language Module generates a set of actions, which are performed by the robots (see Fig. 2 right). In Fig. 3 an example of a set of actions can be observed when all the required ingredients are available. In Fig. 4 the details of a single action are shown.
III-B Large Language Model Module
In this work, we utilized Starling-LM-7B-alpha [23], an open-source LLM developed by UC Berkeley’s NEST Lab using Reinforcement Learning from AI Feedback (RLAIF), as the backbone for the language module, which outperforms GPT-3.5 [24] in several benchmarks, including an 8.09 out of 10 in MT Bench.
III-B1 Semantic planning
Semantic planning involves the use of natural language understanding and generation techniques to interpret human instructions and generate executable plans. In the context of our work, the semantic planner is designed to generate a plan for coordinating the movements of the two robot arm manipulators equipped with different and appropriate tools in order to perform a given action. The planner takes the action description as input and produces a plan specifying the sequence of motions that need to be executed by the robot arms (see Fig. 4). The planner utilizes a template that provides a general structure for the plan. It starts with the “[start of plan]" tag and ends with the “[end of plan]" tag, ensuring that the plan is clearly delineated. The template also includes specific instructions and placeholders that need to be filled based on the task requirements by the vision-language model (VLM) (see section III-C1), as well as a set of rules to follow.
@@ III-B1 Semantic planning
III-B2 Translating semantic plan into APIs function calls with the Code Generator
The plan generated by the semantic planner is translated into a sequence of API call functions. The goal is to convert the textual plan into executable Python code that adheres to a predefined set of functions provided by the API. We use a prompt that outlines the functions available for use, the structure of the plan, and the rules to follow. To illustrate the process, we provide an example plan and its corresponding Python code (Fig. 4), helping to guide the model’s response as in the one-shot prompting technique. This methodology facilitates the seamless translation of high-level task plans into executable Python code, promoting efficiency and clarity.
III-B3 From API function call to Action Generation with The Motion Control APIs
@@
  • open_gripper(manipulator_type): to open the gripper.
  • move_to_object(manipulator_type, object_name): to move the manipulator to the position of an object.
  • grasp(manipulator_type, object_name): to grasp an object.
  • cut(manipulator_type, object_name): for cutting actions.
  • pour(manipulator_type, object_name): for pouring actions.
  • put(manipulator_type, object_name): for pick-and-place actions.
  • toss(manipulator_type, object_name): for mixing objects or vegetables in a bowl.
  • cut_and_put_in(manipulator_type, object_name): which combines the cut and put actions.
The vision language model uses the function get_list_of_objects() to retrieve the names of all objects visible on the table within the camera frame, as well as get_bounding_boxes() which generates the bounding boxes of the desired list of objects. Ensuring explicit naming of API functions is crucial to preventing confusion with the language model. This comprehensive set of API functions provides the necessary control and versatility for robotic operations, enabling a wide range of complex tasks to be performed with precision and efficiency.
III-C Vision Language Model Module
III-C1 Integration of the Qwen Vision-Language Model
The QWEN-VL series are large-scale vision-language models (VLMs) that understand both text and images [25]. Built on the QWEN-LM foundation, they feature visual capabilities through a specialized architecture and training pipeline, achieving state-of-the-art performance across various benchmarks, including image captioning, visual question answering, and visual grounding. In this work, we use the model to verify the availability of required items and provide the 2D-pixel coordinates of these items.
III-C2 Mapping Image Pixel to 3D Coordinates
We assumed that the position of an object is equal to the position of its center. The relationship between the position of an object in the world coordinate system and pixel coordinates is given by the equation:
@@ III-C2 Mapping Image Pixel to 3D Coordinates
${}^{c}_{w}\grave{T}=\left[\begin{array}[]{c|c}{}^{c}_{w}R_{3\times 3}&{}^{c}_{w}t_{3\times 1}\end{array}\right],$ (2)
@@
${}^{c}_{w}T=\left[\begin{array}[]{cccc}&{}^{c}_{w}\grave{T}&&\\ \hline\cr 0&0&0&1\end{array}\right]=\left[\begin{array}[]{c|c}{}^{c}_{w}R_{3\times 3}&{}^{c}_{w}t_{3\times 1}\\ \hline\cr 0_{1\times 3}&1\end{array}\right].$ (3)
@@
Figure 4: Example illustration of Action 2 in Fig. 3. The LLM-based semantic planner receives the action to perform and generates a detailed plan outlining the movements of two robot manipulators. The generated plan is translated into a set of motion API call functions through the LLM-based Code Generator. The execution of the generated code triggers the movement of the two robot manipulators through the Motion Controller.(a)-(b) the manipulator with the gripper moves to grasp the pepper and bring it to the cutting board; (c)-(d) the manipulator with the knife moves to the cutting board, cuts the pepper (e), and places it in the bowl; (f)-(g) the two manipulators return to their initial position.
IV Experimental Evaluation
IV-A Experimental Setup
Two UR3 collaborative robots from Universal Robots were mounted on a Siegmund welding table (see Fig. 1). A custom-designed and 3D-printed camera holder, made of PLA, was attached to the first manipulator, which also held a Logitech HD 1080p webcam and a two-finger Robotiq gripper. A knife holder was similarly designed and installed in the second manipulator’s end effector. Ingredients for various salads were positioned near the first manipulator where the camera was able to capture images to identify the necessary items. A cutting board was placed between the two manipulators for cutting operations, with a plate or bowl in front to place the prepared ingredients. All processes are supervised by a PC (NVIDIA GeForce RTX 4090 24Gb, SSD 2Tb, RAM 64Gb 5600MHz, Intel Core i9, running Ubuntu 22.04).
@@ IV-B Experiments Workflow
As depicted in Fig. 3, we design several custom cooking recipes stored in a vector database, which serves as a set of tasks for the robot manipulators to perform collaboratively. The recipes clearly state the required ingredients as well as the instructions to follow for successful execution. Among these sets of tasks, we evaluate the performance of our system on three selected tasks, including making a vegetable salad, a Russian salad, and a fruit salad. The three tasks are different in nature in the sense that they require several types of actions, such as picking, placing, cutting, pouring, mixing, etc. When a user submits a request, it serves as a query processed through the LLM-Based Retrieval-Augmented Generation (RAG) technique [28], which enhances response quality and factuality by augmenting the language model’s outputs with relevant information from external knowledge sources , in this case, a vector database. After the RAG model produces a response, the VLM module verifies the availability of all the ingredients mentioned in the recipe. Based on the VLM’s response, the LLM module will then generate a list of actions for the two robots to perform if the ingredients are available. Alternatively, if the ingredients are not available or the request is ambiguous, no action will be performed.